# Ihara’s Lemma for $\operatorname{GL}_{d}$: the limit case

### v4

@@ 2.1. Representations of $\operatorname{GL}_{d}(L)$
An irreducible representation $\pi$ of $\operatorname{GL}_{d}(L)$ is said generic, if its restriction to the mirabolic subgroup admits $\tau_{nd}$ as a subquotient.
Let $\pi$ be an irreducible generic ¯$\overline{\mathbb{Q}}_{l}$-representation of $\operatorname{GL}_{d}(L)$ and consider any stable lattice which gives us by modulo $l$ reduction a $\overline{\mathbb{F}}_{l}$- representation uniquely determined up to semi-simplification. Then this modulo $l$ reduction admits an unique generic irreducible constituant.
@@ 2.2. Weil–Deligne inertial types
Recall that a Weil-Deligne representation of $W_{L}$ is a pair $(r,N)$ where
  • $r:W_{L}\longrightarrow\operatorname{GL}(V)$ is a smooth666i.e. continuous for the discrete topology on $V$ representation on a finite dimensional vector space $V$; and
@@
    where $||\bullet||:W_{L}\longrightarrow W_{L}/I_{L}\twoheadrightarrow q^{\mathbb{Z}}$ takes an arithmetic Frobenius element to $q$.
Remark. To a continuous777relatively to the $l$-adic topology on $V$ representation on a finite dimensional ${\mathbb{Q}}_{l}$-vector space $V$, $\rho:W_{L}\longrightarrow\operatorname{GL}(V)$ is attached a Weil-Deligne representation denoted by $\operatorname{WD}(\rho)$. A Weil representation of $W_{L}$ is also said of Galois type if it comes from a representation of $G_{L}$.
Main example: let $\rho:W_{L}\longrightarrow\operatorname{GL}(V)$ be a smooth irreducible representation on a finite dimensional vector space $V$. For $k\geq 1$ an integer, we then define a Weil-Deligne representation
@@
where for $20\leq i\leq k-$ 2, the isomorphism $N:V(i)\cong V(i+1)$ is induced by some choice of a basis of $\overline{L}(1)$ and $N_{|V(k-1)}$ is zero. Then every Frobenius semi-simple Weil-Deligne representation of $W_{L}$ is isomorphic to some $\bigoplus_{i=1}^{r}\operatorname{Sp}(\rho_{i},k_{i})$, for smooth irreducible representations $\rho_{i}:W_{L}\longrightarrow\operatorname{GL}(V_{i})$ and integers $k_{i}\geq 1$. Up to obvious reorderings, such a writing is unique.
Let now $\rho$ be a continuous representation of $W_{L}$, or its Weil-Deligne representation $\operatorname{WD}(\rho)$, and consider its restriction to $I_{L}$, $\tau:=\rho_{|I_{L}}$. Such an isomophism class of a finite dimensional continuous representation of $I_{L}$ is then called an inertial type.
@@
Let $\mathcal{I}_{0}$ the set of inertial types that extend to a continuous irreducible representation of $G_{L}$.
@@
Let $\operatorname{Part}$ be the set of decreasing sequences of positive integers $\underline{d}=(\underline{d}(1)\geq\underline{d}(2)\geq\cdots)$ viewed as a partition of $\sum\underline{d}:=\sum_{i}\underline{d}(i)$.
@@
Let $f:\mathcal{I}_{0}\longrightarrow\operatorname{Part}$ with finite support. We then denote by $\tau_{f}$ the restriction to $I_{L}$ of
@@
where $\rho_{\tau_{0}}$ is a fixed extension of $\tau_{0}$ to $W_{L}$.
@@
We let $\mathrm{rec}_{L}$ denote the local reciprocity map of [HT01, Theorem A]. Fix an isomorphism ı¯superscriptsimilar-toitalic-$\imath\overline{{\mathbb{Q}}}_{\ell}\buildrel\sim\over{\rightarrow}{\mathbb{C}}$. We normalize the local reciprocity map $\mathrm{rec}$ of [HT01, Theorem A], defined on isomorphism classes of irreducible smooth representations of $\operatorname{GL}_{n}(L)$ over ${\mathbb{C}}$ as follows: if $\pi$ is the isomorphism class of an irreducible smooth representation of $\operatorname{GL}_{n}(L)$ over $\overline{{\mathbb{Q}}}_{\ell}$, then
$\rho_{\ell}(\pi)\stackrel{{\scriptstyle\textrm{\tiny{{def}}}}}{{=}}\imath^{-1}\circ\mathrm{rec}_{L}\circ\imath(\pi\otimes_{\overline{{\mathbb{Q}}}_{\ell}}|\det|^{(1-n)/2}).$
Then $\rho_{\ell}(\pi)$ is the isomorphism class of an $n$-dimensional, Frobenius semisimple Weil–Deligne representation of $W_{L}$ over $\overline{{\mathbb{Q}}}_{\ell}$, independent of the choice of $\imath$. Moreover, if $\rho$ is an isomorphism class of an $n$-dimensional, Frobenius semisimple Weil–Deligne representation of $W_{L}$ over $M$, then $\rho_{\ell}^{-1}(\rho)$ is defined over $M$ (cf. [CEG+16, §1.8]).
@@
If $\pi$ and $\pi^{\prime}$ are irreducible generic representations of $\operatorname{GL}_{d}(L)$ such that $\rho_{\ell}(\pi)|I_{L}\cong\rho_{\ell}(\pi^{\prime})|I_{L}$ then $\pi_{|\operatorname{GL}_{d}(\mathcal{O}_{L})}\cong\pi^{\prime}_{|\operatorname{GL}_{d}(\mathcal{O}_{L})}.$

### v5

@@ 2.1. Representations of $\operatorname{GL}_{d}(M)$
An irreducible representation $\pi$ of $\operatorname{GL}_{d}(M)$ is said generic, if its restriction to the mirabolic subgroup admits $\tau_{nd}$ as a subquotient.
Let $\pi$ be an irreducible generic ¯$\overline{\mathbb{Q}}_{l}$-representation of $\operatorname{GL}_{d}(M)$ and consider any stable lattice which gives us by modulo $l$ reduction a $\overline{\mathbb{F}}_{l}$- representation uniquely determined up to semi-simplification. Then this modulo $l$ reduction admits an unique generic irreducible constituant.
Lemma 2.1.6.
Let $1\leq t\leq s$ and consider
  • $\underline{s}=(l_{1}\geq\cdots\geq l_{r})$ a partition of $s$ with $l_{1}<t$;
  • $\pi_{1},\cdots,\pi_{r}$ irreducible cuspidal representations whose modulo $l$ reductions are isomorphic to $\varrho$.
There exists an irreducible subquotient $\tau$ of $\mathrm{st}_{l_{1}}(\pi_{1})\times\cdots\times\mathrm{st}_{l_{r}}(\pi_{r})$ such that
  • whatever is $\underline{s}^{\prime}=(l^{\prime}_{1}\geq\cdots\geq l^{\prime}_{r^{\prime}})$ with $l^{\prime}_{1}\geq t$,
  • and $\pi^{\prime}_{1},\cdots,\pi^{\prime}_{r^{\prime}}$ irreducible cuspidal representations with modulo $l$ reduction isomorphic to $\varrho$,
then $\tau$ is not a subquotient of the modulo $l$ reduction of $\mathrm{st}_{l^{\prime}_{1}}(\pi^{\prime}_{1})\times\cdots\times\mathrm{st}_{l^{\prime}_{r}}(\pi^{\prime}_{r^{\prime}})$.
Proof.
Note that $\mathrm{st}_{l_{1}}(\pi_{1})\times\cdots\times\mathrm{st}_{l_{r}}(\pi_{r})$ has the same modulo $l$ reduction as
$\mathrm{st}_{l_{1}}(\pi_{1}\{\frac{l_{1}-s}{2}\})\times\mathrm{st}_{l_{2}}(\pi_{2}\{\frac{l_{2}-s}{2}+l_{1}-1\}\times\cdots\times\mathrm{st}_{l_{r}}(\pi_{1}\{\frac{l_{r}-s}{2}+l_{1}+\cdots+l_{r-1}\})$ (2)
where the shifts are chosen so that $\mathrm{st}_{l_{1}}(\pi\{\frac{l_{1}-s}{2}\})\times\cdots\times\mathrm{st}_{l_{r}}(\pi\{\frac{s-l_{r}}{2}\})$ is the subquotient of (1) associated to
$(\overbrace{1,\cdots,1}^{l_{1}-1},2,\overbrace{1,\cdots,1}^{l_{1}-1},2,\cdots,\overbrace{1,\cdots,1}^{l_{r}-1})\in\operatorname{Dec}(s).$
Note that this irreducible constituant of (1) is the less non degenerate subquotient and we denote by $\tau$ is its modulo $l$ reduction which remains irreducible.
The property stated in the lemma then follows from the fact that whatever is an irreducible subquotient of the modulo $l$ reduction of $\mathrm{st}_{l^{\prime}_{1}}(\pi^{\prime}_{1})\times\cdots\times\mathrm{st}_{l^{\prime}_{r}}(\pi^{\prime}_{r^{\prime}})$, it has a non zero derivative of order $l^{\prime}_{1}$, and so a non zero derivative of order $t$, while the derivative of order $t$ of $\tau$ is zero.
Definition 2.1.7.
Elements $\tau$ constructed in the above lemma will be said to be $(\varrho,t)$-small.
Consider any non degenerate irreducible representation $\Pi:=\mathrm{st}_{l_{1}}(\pi_{1})\times\cdots\times\mathrm{st}_{l_{r}}(\pi_{r})$ where the modulo $l$ reduction of $\pi_{1},\cdots,\pi_{r}$ is isomorphic to $\varrho$. As the modulo $l$ reduction of $\Pi$ contains an unique irreducible non degenerate subquotient, there exists then an unique stable lattice such that its modulo $l$ reduction has an irreducible generic socle $\tau_{gen}$, cf. [EGS15] lemma 4.1.1.
Notation 2.1.8.
We denote by $\Gamma(\Pi)^{gen}$ the lattice of $\Pi$ for which the socle of its modulo $l$ reduction is generic.
@@ 2.2. Weil–Deligne inertial types
Recall that a Weil-Deligne representation of $W_{M}$ is a pair $(r,N)$ where
  • $r:W_{M}\longrightarrow\operatorname{GL}(V)$ is a smooth777i.e. continuous for the discrete topology on $V$ representation on a finite dimensional vector space $V$; and
@@
    where $||\bullet||:W_{M}\rightarrow W_{M}/I_{M}\twoheadrightarrow q^{\mathbb{Z}}$ takes an arithmetic Frobenius element to $q$.
Remark. To a continuous888relatively to the $l$-adic topology on $V$ representation on a finite dimensional ${\mathbb{Q}}_{l}$-vector space $V$, $\rho:W_{M}\longrightarrow\operatorname{GL}(V)$ is attached a Weil-Deligne representation denoted by $\operatorname{WD}(\rho)$. A Weil representation of $W_{M}$ is also said of Galois type if it comes from a representation of $G_{M}$.
Main example: let $\rho:W_{M}\longrightarrow\operatorname{GL}(V)$ be a smooth irreducible representation on a finite dimensional vector space $V$. For $k\geq 1$ an integer, we then define a Weil-Deligne representation
@@
where for $20\leq i\leq k-$ 2, the isomorphism $N:V(i)\cong V(i+1)$ is induced by some choice of a basis of $\overline{L}(1)$ and $N_{|V(k-1)}$ is zero. Then every Frobenius semi-simple Weil-Deligne representation of $W_{M}$ is isomorphic to some $\bigoplus_{i=1}^{r}\operatorname{Sp}(\rho_{i},k_{i})$, for smooth irreducible representations $\rho_{i}:W_{M}\longrightarrow\operatorname{GL}(V_{i})$ and integers $k_{i}\geq 1$. Up to obvious reorderings, such a writing is unique.
Let now $\rho$ be a continuous representation of $W_{M}$, or its Weil-Deligne representation $\operatorname{WD}(\rho)$, and consider its restriction to $I_{M}$, $\tau:=\rho_{|I_{M}}$. Such an isomophism class of a finite dimensional continuous representation of $I_{M}$ is then called an inertial type.
@@
Let $\mathcal{I}_{0}$ the set of inertial types that extend to a continuous irreducible representation of $G_{M}$.
@@
Let $\operatorname{Part}$ be the set of decreasing sequences of positive integers
$\underline{d}=(\underline{d}(1)\geq\underline{d}(2)\geq\cdots)$
viewed as a partition of $\sum\underline{d}:=\sum_{i}\underline{d}(i)$. We also denote by $\operatorname{Part}(s)$ the set of partition of $s$: $\operatorname{Part}=\coprod_{s\geq 1}\operatorname{Part}(s)$.
@@
Let $f:\mathcal{I}_{0}\longrightarrow\operatorname{Part}$ with finite support. We then denote by $\tau_{f}$ the restriction to $I_{M}$ of
@@
where $\rho_{\tau_{0}}$ is a fixed extension of $\tau_{0}$ to $W_{M}$.
@@
We let $\mathrm{rec}_{M}$ denote the local reciprocity map of [HT01, Theorem A]. Fix an isomorphism ı:¯:italic-$\imath:\overline{{\mathbb{Q}}}_{\ell}\buildrel\sim\over{\rightarrow}{\mathbb{C}}$. We normalize the local reciprocity map $\mathrm{rec}$ of [HT01, Theorem A], defined on isomorphism classes of irreducible smooth representations of $\operatorname{GL}_{n}(M)$ over ${\mathbb{C}}$ as follows: if $\pi$ is the isomorphism class of an irreducible smooth representation of $\operatorname{GL}_{n}(M)$ over $\overline{{\mathbb{Q}}}_{\ell}$, then
$\rho_{\ell}(\pi)\stackrel{{\scriptstyle\textrm{\tiny{{def}}}}}{{=}}\imath^{-1}\circ\mathrm{rec}_{M}\circ\imath(\pi\otimes_{\overline{{\mathbb{Q}}}_{\ell}}|\det|^{(1-n)/2}).$
Then $\rho_{\ell}(\pi)$ is the isomorphism class of an $n$-dimensional, Frobenius semisimple Weil–Deligne representation of $W_{M}$ over $\overline{{\mathbb{Q}}}_{\ell}$, independent of the choice of $\imath$. Moreover, if $\rho$ is an isomorphism class of an $n$-dimensional, Frobenius semisimple Weil–Deligne representation of $W_{M}$ over $M$, then $\rho_{\ell}^{-1}(\rho)$ is defined over $M$ (cf. [CEG+16, §1.8]).
@@
If $\pi$ and $\pi^{\prime}$ are irreducible generic representations of $\operatorname{GL}_{d}(M)$ such that $\rho_{\ell}(\pi)|I_{M}\cong\rho_{\ell}(\pi^{\prime})|I_{M}$ then $\pi_{|\operatorname{GL}_{d}(\mathcal{O}_{M})}\cong\pi^{\prime}_{|\operatorname{GL}_{d}(\mathcal{O}_{M})}.$
Theorem 2.2.4.
(cf. [BC09] proposition 6.3.3 or [Sho18] theorem 3.7)
Let $G=\operatorname{GL}_{n}$. Let $\tau$ be a inertial type for $M$. Then there is a smooth irreducible $\operatorname{GL}_{n}({\mathcal{O}}_{M})$-representation $\sigma(\tau)$ over $E$ such that for any irreducible admissible representation $\pi$ of $\operatorname{GL}_{n}(M)$, one has:
  1. ((i))
    if $\pi|_{\operatorname{GL}_{n}({\mathcal{O}}_{M})}$ contains $\sigma(\tau)$ then $\rho_{\ell}(\pi)|_{I_{M}}\preceq\tau$;
  2. ((ii))
    if $\rho_{\ell}(\pi)|_{I_{M}}=\tau$, then $\pi|_{\operatorname{GL}_{n}({\mathcal{O}}_{M})}$ contains $\sigma(\tau)$ with multiplicity one;
  3. ((iii))
    if $\rho_{\ell}(\pi)|_{I_{M}}\preceq\tau$ and $\pi$ is generic, then $\pi|_{\operatorname{GL}_{n}({\mathcal{O}}_{M})}$ contains $\sigma(\tau)$ and the multiplicity is one if furthermore $\tau$ is maximal with respect to $\preceq$.
Remark. For example take $\tau_{0}$ the trivial representation and consider the following partitions $\underline{d}=(1\geq 1\geq\cdots\geq 1)$ (resp. $(d)$) denoted also by $(1^{d})$. Denote then by $\tau$ the associated inertial type of notation 2.2.2. Then $\sigma(\tau)$ is the trivial representation (resp. is inflated from the Steinberg representation of $\operatorname{GL}_{d}(\kappa)$). Note also that $\pi$ contains $\sigma(\tau)$ if and only if $\pi$ is unramified (resp. it implies that $r_{\ell}(\pi)|I_{M}$ is unipotent).
We need more details about the construction of $\sigma(\tau)$ which rests on the notion of SZ-stratum of [SZ99]. We first recall quickly the basic notions of type theory of Bushnell and Kutzko. Let ${\mathbb{K}}$ denote ¯$\overline{\mathbb{Q}}_{l}$ or $\overline{\mathbb{F}}_{l}$. Let $V$ be a vector space over $F$ and let $G=\operatorname{Aut}_{F}(V)$ and $A=\operatorname{End}_{F}(V)$.
[…]
- The hereditary ${\mathcal{O}}_{M}$-orders in $A$ are those orders ${\mathfrak{U}}$ that arise as the stabiliser of some ${\mathcal{O}}_{F}$-lattice chain; it is maximal if and only if it stabilises a lattice chain of period $e=1$.
- A hereditary order ${\mathfrak{U}}\subseteq A$ has a unique two-sided maximal ideal $\mathfrak{P}=\{x\in{\mathfrak{U}}:x\Lambda_{i}\subseteq\Lambda_{i+1}\hbox{ forall }i\in{\mathbb{Z}}\}$. We write $U({\mathfrak{U}})$ for the group of units in ${\mathfrak{U}}$ and $U^{1}({\mathfrak{U}}):=1+\mathfrak{P}$.
Definition 2.2.5.
A simple strata in $A$ is $[{\mathfrak{U}},m,0,\beta]$ where
  • ${\mathfrak{U}}$ is a hereditary ${\mathcal{O}}_{K}$-order in $A$;
  • $m>0$ is an integer;
  • $\beta\in\mathfrak{P}^{-m}\setminus\mathfrak{P}^{1-m}$ is such that $E:=L[\beta]$ is a field and $E^{\times}$ is contained in the normalisez of $U({\mathfrak{U}})$;
  • the number $k_{0}(\beta,{\mathfrak{U}}(E))$ defined in [BK93] §1.4 is strictly negative.
The chain lattice defining ${\mathfrak{U}}$ can be seen as an ${\mathcal{O}}_{E}$-lattice chain and we denote ${\mathfrak{B}}:={\mathfrak{U}}\cap\operatorname{End}_{E}(V)$ its stabiliser. We define the groups $U({\mathfrak{B}})$ and $U^{1}({\mathfrak{B}})$ as for ${\mathfrak{U}}$. In [BK93], the authors associate to the simple stratum $[{\mathfrak{U}},m,0,\beta]$, compact open subgroups $J=J(\beta,{\mathfrak{U}})$, $J^{1}=J^{1}(\beta,{\mathfrak{U}})$ and $H=H^{1}(\beta,{\mathfrak{U}})$ of $U({\mathfrak{U}})$ such that:
  • $J^{1}$ is a normal prop-$p$- subgroup of $J$;
  • $H^{1}$ is a normal subgroup of $J^{1}$;
  • $U({\mathfrak{B}})\subseteq J$ and $U^{1}({\mathfrak{B}})\subseteq J^{1}$ and the induced map
    $U({\mathfrak{B}})/U^{1}({\mathfrak{B}})\longrightarrow J/J^{1}$
    is a isomorphism.
By [BK93] §3.2, there is a set ${\mathcal{C}}({\mathfrak{U}},0,\beta)$ of simple characters of $H^{1}(\beta,{\mathfrak{U}})$. For $\theta\in{\mathcal{C}}({\mathfrak{U}},0,\beta)$ there is a unique irreducible representation $\eta$ of $J^{1}(\beta,{\mathfrak{U}})$ whose restriction to $H^{1}(\beta,{\mathfrak{U}})$ contains $\theta$. There is then a distinguished class of $\beta$-extensions $\kappa$ of $\eta$ to $J(\beta,{\mathfrak{U}})$.
It is the main result of [BK99] that every Berstein component of $\operatorname{Rep}_{\mathbb{K}}(G)$ has a type and give an explicit construction of them. When $\Omega$ is supercuspidal then they construct a type $(J,\lambda)$ such that $J=J(\beta,{\mathfrak{U}})$ for a simple stratum $[{\mathfrak{U}},m,0,\beta]$ in $A$ in which ${\mathfrak{B}}$ is a maximal ${\mathcal{O}}_{E}$-order and $\lambda=\kappa\otimes\nu$ where $\kappa$ is a $\beta$-extension of an irreducible representation $\eta$ containing a simple character $\theta\in{\mathcal{C}}({\mathfrak{U}},0,\beta)$ and $\nu$ is a cuspidal representation of $J/J^{1}\cong\operatorname{GL}_{n/[E:L]}(k_{E})$. This type is called maximal.
From [BK99] §4, the character $\theta$ determines a ps-character $(\Theta,0,\beta)$ viewed as a function $\Theta$ on the set of simple strata $[{\mathfrak{U}},m,0,\beta]$ taking such a stratum to an element $\Theta({\mathfrak{U}})\in{\mathcal{C}}({\mathfrak{U}},0,\beta)$. By [BK99] §4.5, the endo-class of this ps-character is determine by $\Omega$.
If $[M,\pi]$ is an inertial equivalence class of supercuspidal pair corresponding to a Berstein component $\Omega$ of $\operatorname{Rep}_{\mathbb{K}}(G)$, then let $(J_{M},\lambda_{M})$ be a maximal type for the supercuspidal Berstein component $\Omega_{M}$ of $\operatorname{Rep}_{\mathbb{K}}(M)$ containing $\pi$. By the results of [BK99], there is, what is called a $G$-cover, $(J,\lambda)$ of $(J_{M},\lambda_{M})$ which is a type for $\Omega$.
Definition 2.2.6.
(cf. [Sho18] definition 6.11)
A SZ-datum over ${\mathbb{K}}$, is a set
$\mathfrak{S}=\Bigl{\{}(E_{i},\beta_{i},V_{i},{\mathfrak{B}}_{i},\lambda_{i}),~{}i=1,\cdots,r\Bigr{\}}$
where $r$ is a positive integer and for each $i=1,\cdots,r$ we have
  • $E_{i}/F$ is a finite extension generated by an element $\beta_{i}\in E_{i}$,
  • $V_{i}$ is an $E_{i}$-vector space of finite dimension $N_{i}$,
  • ${\mathfrak{B}}_{i}\subseteq\operatorname{End}_{E_{i}}(V_{i})$ is a maximal hereditary ${\mathcal{O}}_{E_{i}}$-order,
  • let denote ${\mathfrak{U}}_{i}$ the associated ${\mathcal{O}}_{M}$-order in $A_{i}:=\operatorname{End}_{M}(V_{i})$ and let $m_{i}:=-v_{E_{i}}(\beta_{i})$. Then $[{\mathfrak{U}}_{i},m_{i},0,\beta_{i}]$ is a simple stratum and $\lambda_{i}$ is an ${\mathbb{K}}$-representation of $J_{i}:=J(\beta_{i},{\mathfrak{U}}_{i})$ of the form $\kappa_{i}\otimes\nu_{i}$ where $\kappa_{i}$ is a $\beta_{i}$-extension of the representation $\eta_{i}$ of $J^{1}_{i}:=J^{1}(\beta_{i},{\mathfrak{U}}_{i})$ containing some simple character $\theta_{i}\in{\mathcal{C}}({\mathfrak{U}}_{i},0,\beta_{i})$ of $H^{1}_{i}=H^{1}(\beta_{i},{\mathfrak{U}}_{i})$ and $\nu_{i}$ is an irreducible representation of $U({\mathfrak{B}}_{i})/U^{1}({\mathfrak{B}}_{i})\cong\operatorname{GL}_{N_{i}}(k_{E_{i}})$ over ${\mathbb{K}}$;
  • no two of the $\theta_{i}$ are endo-equivalent in the sense of [BK99] §4.
Let $V=\bigoplus_{i=1}^{r}V_{i}$, $A=\operatorname{End}_{M}(V)$ and $G=\operatorname{Aut}_{M}(V)$. The Levi subgroup $M=\prod_{i=1}^{r}\operatorname{Aut}_{M}(A_{i})\subseteq G$ has compact open subgroups $J^{1}_{M}\triangleleft J_{M}$ where $J^{1}_{M}=\prod_{i=1}^{r}J^{1}_{i}$ and similarly for $J_{M}$. Let denote by $\eta_{M}=\bigotimes_{i=1}^{r}\eta_{i}$, a representation of $J^{1}_{M}$ and similarly for the representations $\kappa_{M}$ and $\lambda_{M}$ of $J_{M}$. Note that $\eta_{M}$ and $\kappa_{M}$ are clearly irreducible and $\lambda_{M}$ is irreducible by [Sho18] propositions 6.12.
Since no two of the $\theta_{i}$ are endo-equivalent, by [BK99] §8, see also [MS14] proposition 2.28, we have compact open subgroups $J$ and $J^{1}$ of $G$ and representations $\eta$ of $J^{1}$, $\kappa$ of $J$ and $\lambda$ of $J$ such that $(J^{1},\eta)$ (resp. $(J,\kappa)$, resp. $(J,\lambda)$) is a $G$-cover of $(J^{1}_{M},\eta_{M})$ (resp. $(J_{M},\kappa_{M})$, resp. $(J_{M},\lambda_{M})$) and $J/J^{1}=J_{M}/J^{1}_{M}$ with $\lambda=\kappa\otimes(\bigotimes_{i=1}^{r}\nu_{i})$ under this identification.
Definition 2.2.7.
For $\mathfrak{S}$ an SZ-stratum and $J$ and $\lambda$ as above. Let $K$ be a maximal compact subgroup of $G$ such that $J_{M}\subseteq K\cap M$. We then denote by
$\sigma(\mathfrak{S}):=\operatorname{ind}_{J}^{K}(\lambda),$
which by [Sho18] theorem 6.16 is irreducible.
Start now from ${\mathcal{P}}\in{\mathcal{I}}=\{f:{\mathcal{I}}_{0}\longrightarrow\operatorname{Part}\}$ and let $n=\sum_{\tau_{0}\in{\mathcal{I}}_{0}}\sum f(\tau_{0})\dim\tau_{0}$. Let $V$ be a $L$-vector space of dimension $N$ and let $G=\operatorname{Aut}_{M}(V)$. Let $(M^{0},\pi)$ be a supercuspidal pair in the inertial equivalence class associated to the Bernstein component attached to ${\mathcal{P}}$. Write $M^{0}=\prod_{i=1}^{t}M_{i}^{0}$ where each $M_{i}^{0}$ is the stabiliser of some $n_{i}$-dimensional subspace $V_{i}^{0}$ of $V$. Write $\pi=\bigotimes_{i=1}^{t}\pi_{i}$ and let $\Omega_{i}$ be the supercuspidal Berstein component containing $\pi_{i}$. For each $\Omega_{i}$ there is an associated endo-class of ps-character $\Theta_{i}^{0}=(\Theta_{i}^{0},0,\beta_{i}^{0})$. Let $M=\prod_{i=1}^{r}M_{i}$ the Levi subgroup of $G$ obtaining from $M$ by gathering $M_{j}^{0}$ with $M_{k}^{0}$ if and only if $\Theta_{j}^{0}=\Theta_{k}^{0}$: we then write simply $(\Theta_{i},0,\beta_{i})$ for the common value.
We attach to ${\mathcal{P}}$ a SZ-datum
$\mathfrak{S}_{\mathcal{P}}=\Bigl{\{}(E_{i},\beta_{i},V_{i},{\mathfrak{B}}_{i},\kappa_{i}\otimes\nu_{i}):i=1,\cdots,r\Bigr{\}}$
as follows. Suppose first that $r=1$ and wrtie $(\Theta,0,\beta)$ for the common value of $(\Theta_{i}^{0},0,\beta_{i}^{0})$.
  • $E=L[\beta]$;
  • let $(J_{i}^{0},\lambda_{i}^{0})$ be the maximal simple type for the supercuspidal Berstein component $\Omega_{i}^{0}$ with $J_{i}^{0}=J(\beta,UF_{i}^{0})$ for a simple stratum $[{\mathfrak{U}}_{i}^{0},m_{i}^{0},0,\beta_{i}^{0}]$ and $\lambda_{i}^{0}$ contains $\theta_{i}^{0}:=\Theta({\mathfrak{U}}_{i}^{0},0,\beta_{i})$. As above we then have compact open subgroups $J^{1}\subseteq J$ of $G$ and a representation $\eta$ of $J^{1}$ containing the simple character $\Theta({\mathfrak{U}},0,\beta)$ where ${\mathfrak{U}}$ is a hereditary ${\mathcal{O}}_{M}$-order in $A$ and ${\mathcal{U}}_{i}\cap B={\mathfrak{B}}$ is a maximal hereditary ${\mathcal{O}}_{E}$-order.
  • We choose compatible $\beta$-extensions, in the sense of [Sho18] §6.6, $\kappa_{i}^{0}$ of $\eta_{i}^{0}$ coming from a $\beta$-extension $\kappa$ of $\eta$, and decompose each $\lambda_{i}^{0}$ as $\kappa_{i}^{0}\otimes\nu_{i}^{0}$ where $\nu_{i}^{0}$ is a cuspidal representation of $J_{i}^{0}/J_{i}^{1,0}=U({\mathfrak{B}}_{i}^{0})/U^{1}({\mathfrak{B}}_{i}^{0})\cong\operatorname{GL}_{n_{i}^{0}/[E:L]}(k_{E})$ for an integer $n_{i}^{0}$. Then $J/J^{1}\cong\operatorname{GL}_{n/[E:L]}(k_{E})$ so that we can view each $\nu_{i}^{0}$ as an element of ¯$\overline{\mathcal{I}}_{0}$ and define an element $\overline{\mathcal{P}}\in\overline{\mathcal{I}}$ by $\overline{\mathcal{P}}(\nu_{i}^{0})={\mathcal{P}}(\tau_{i})$ where $\tau_{i}\in{\mathcal{I}}_{0}$ corresponds to $\Omega_{i}$.
  • We then write $\nu=\sigma_{\overline{\mathcal{P}}}$ a representation of $\operatorname{GL}_{n/[E:L]}(k_{E})$ and see it as a representation of $J/J^{1}$.
  • We repeat this construction for every $i=1,\cdots,r$.
Notation 2.2.8.
For $\tau=\tau_{\mathcal{P}}$ we write $\sigma(\tau):=\sigma(\mathfrak{S}_{\mathcal{P}})$.
Theorem 2.2.9.
([Sho18] 6.20)
Let ${\mathcal{P}}^{\prime}\in{\mathcal{I}}$ with degree $n$ and let $(M^{\prime},\pi^{\prime})$ be any discrete pair in the inertial equivalence class associated to ${\mathcal{P}}^{\prime}$. For any parabolic subgroup $Q^{\prime}$ of $G$ with Levi $M^{\prime}$, we have
$\dim\operatorname{Hom}_{K}(\sigma(\mathfrak{S}_{\mathcal{P}}),\operatorname{ind}_{Q^{\prime}}^{G}\pi^{\prime})=\prod_{\tau_{0}\in{\mathcal{I}}_{0}}m({\mathcal{P}}(\tau_{0}),{\mathcal{P}}^{\prime}(\tau_{0})),$
where for any two partition $\underline{\lambda},\underline{\mu}$ of $n$, $m(\underline{\lambda},\underline{\mu})$ is the usual Kostka number which counts the number of possible ways to fill the Young tableau with lines of respective size $\lambda_{1},\cdots$ with $\mu_{1}$ one, $\mu_{2}$ two and so on, with the following properties: the sequence of labels in each columns is strictly increasing, while it is increasing in each lines.
Note that
  • if $\underline{\lambda}<\underline{\mu}$ then $m(\underline{\lambda},\underline{\mu})=0$;
  • $m(\underline{\lambda},\underline{\lambda})=1$;
  • $m((n),\underline{\mu})=1$.
In particular if ${\mathcal{P}}$ is maximal that is if for every $\tau_{0}\in{\mathcal{I}}_{0}$ we have ${\mathcal{P}}(\tau_{0})=(\sum{\mathcal{P}}(\tau_{0}))$, then the multiplicity of $\sigma(\mathfrak{S}_{\mathcal{P}})$ in $\operatorname{ind}_{Q^{\prime}}^{G}\pi^{\prime}$ is $1$ if $\sum{\mathcal{P}}^{\prime}(\tau_{0})=\sum{\mathcal{P}}(\tau_{0})$ for every $\tau_{0}\in{\mathcal{I}}_{0}$, otherwise it is 00.
Let $\mathfrak{S}=\Bigl{\{}(E_{i},\beta_{i},V_{i},{\mathfrak{B}}_{i},\lambda_{i});i=1,\cdots,r\Bigr{\}}$ be an SZ-datum. Write $\lambda_{i}=\kappa_{i}\otimes\nu_{i}$ where the $\nu_{i}$ are irreducible representations of $J_{i}/J^{1}_{i}$. Note that the modulo $l$ reduction $\overline{\kappa}_{i}$ of $\kappa_{i}$ remains irreducible. We then decompose the semisimplification of the modulo $l$ reduction of $\nu_{i}$:
$\overline{\nu}_{i}^{ss}=\bigoplus_{j\in S_{i}}\mu_{i,j}\nu_{i,j}$
where $S_{i}$ is some finite indexing set, $\nu_{i,j}$ are distinct irreducible $\overline{\mathbb{F}}_{l}$-representations and $\mu_{i,j}\in{\mathbb{N}}$. For $\underline{j}=(j_{1},\cdots,j_{r})\in S_{1}\times\cdots\times S_{r}$, define an SZ-datum $\mathfrak{S}_{\underline{j}}$ over $\overline{\mathbb{F}}_{l}$ by
$\mathfrak{S}_{\underline{j}}=\Bigl{\{}(E_{i},\beta_{i},V_{i},{\mathfrak{B}}_{i},\overline{\kappa}_{i}\otimes\nu_{i,j_{i}});i=1,\cdots,r\Bigr{\}}.$
Theorem 2.2.10.
(cf [Sho18] 6.23)
The semisimplified modulo $l$ reduction of $\sigma(\mathfrak{S})$ is
$\bigoplus_{\underline{j}\in S_{1}\times\cdots\times S_{r}}\mu_{\underline{j}}\sigma(\mathfrak{S}_{\underline{j}}),$
where $\mu_{\underline{j}}:=\prod_{i=1}^{r}\mu_{i,j_{i}}$.